# Shift-cyclicity in Analytic Function Spaces

### v1

@@ 1. SCP
    Note that $f(z)=z-w$ satisfies (2.2) exactly when $w\not\in\mathbb{D}$ [[, see]Lemma 15.17]Rud74. If we combine Theorem 2.10 and Proposition 1.12, we get that $P$ is shift-cyclic in any of the Hardy spaces if and only if it is non-vanishing on $\mathbb{D}$. For a proof that does not use Theorem 2.10, see [GNN70, Theorem 5].
@@
    For any $f\in H^{p}(\mathbb{D})$ with $0<p<\infty$, a result of Garnett [Gar06, Chapter II Theorem 7.4] then shows that
@@ 5. Inv
SCP refers to the shift-cyclicity of polynomials. It will be a general theme moving forward that the vanishing set of a polynomial plays crucial role in determining its shift-cyclicity.
@@
Multiplier-cyclicity is a natural substitute for shift-cyclicity in non-analytic function spaces. We learn more about this topic in Section 4.2.
@@ 2.2. On the unit polydisk $\mathbb{D}^{d}$
The boundary properties of $H^{p}(\mathbb{D})$ functions are explained by the results of Fatou [Fat06] and Riesz [Rie23]. These results are, in essence, consequences of the symmetries of the Poisson kernel and the Poisson integral representation of Hardy space functions. Many of these properties can be generalized to $d>1$ via slice functions, i.e., given any $f\in\operatorname{Hol}(\mathbb{D}^{d})$ and $\zeta\in\mathbb{T}^{d}$, consider
@@
exist a.e. on $\mathbb{T}^{d}$ [[, see]Theorem 3.3.3]Rud69. Bochner [Boc69] used slice functions to extend Riesz’s theorem and showed that $\log|f^{*}|\in L^{1}(\mathbb{T}^{d})$ unless $f\equiv 0$, and
@@
for all $f\in H^{p}(\mathbb{D}^{d})$ with $0<p<\infty$ (see [Rud69, Theorem 3.4.3] for a different treatment). It is then not difficult to verify that Corollary 2.6 and Theorem 2.7 can be generalized for $d>1$. Unfortunately, we cannot go too far from here in terms of factorization and shift-cyclicity when $d>1$.
@@ 2.2.1. Negative results
All examples mentioned in this subsection are non-trivial, as well as technical, so we omit the details to keep our discussion compact. The reader is encouraged to look at the references below for these details. Let us start with the following result of Rudin [[, see]corollary to Theorem 4.4.2]Rud69 which shows that Theorem 2.10 cannot be generalized as is to the case $d>1$.
@@
One might wonder if we could provide an appropriate description of those shift-invariant subspaces that are finitely generated. The following result of Miles [Mil73] sheds some light on this. Similar to the one variable case, we let $\mathcal{Z}(f)$ denote the zero set of any $f\in\operatorname{Hol}(\mathbb{D}^{d})$ in $\mathbb{D}^{d}$ counting multiplicities [^5] The multiplicity of a zero at $w$ is the smallest integer $k$ such that the $k$-th term in the homogeneous expansion of $f$ around $w$ does not vanish at $w$ [[, see]Section 1.1]Rud69..
@@
Combined with Theorem 2.12, the above equation gives the following result. Note that while Miles’ result holds when $d=2$, one can easily generalize it to any $d>1$ by taking
@@
The penultimate nail in the coffin is that there is no real hope for a useful factorization result for $d>1$ either. Firstly, Theorem 2.12 shows that the result of F. and R. Nevanlinna [NN22] (see p. 2.3) does not hold for $d>1$. Next, we have the following result of Rosay [Ros75] (generalizing a result of Rudin [Rud66] and Miles [Mil75]). Recall from Mult that the Smirnov factorization result implies that every $H^{1}(\mathbb{D})$ function can be factored into a product of two $H^{2}(\mathbb{D})$ functions.
@@
Finally, if one only cares about shift-cyclicity of a function, then it is natural to ask if there is an analogue of an outer function when $d>1$. Rudin defines an outer function as a function $f\in H^{p}(\mathbb{D}^{d})$ such that
$-\infty<\log|f(0)|=\int_{\mathbb{T}^{d}}\log|f^{*}|,$
@@
In [Rud69, Theorem 4.4.6], Rudin shows that every shift-cyclic function in $H^{p}(\mathbb{D}^{d})$ is outer. Thus, one-half of the Smirnov–Beurling holds. However, in [Rud69, Theorem 4.4.8 (b)], Rudin then constructs an explicit example of an outer function in $H^{2}(\mathbb{D}^{2})$ that is not shift-cyclic.
@@ 2.2.2. Positive results
We start with a result of Ahern and Clark which shows that all shift-invariant subspaces of $H^{2}(\mathbb{D}^{d})$ with finite codimension are generated by finitely many polynomials [[, see]Theorem 3 and Corollary 4]AC70.
@@
Finite codimensionality of polynomial ideals is fairly well-understood, both algebraically and computationally [[, see]Section 3.7]KR00. In fact, the codimension of such an ideal is precisely the number of common zeros (in $\mathbb{D}^{d}$) of polynomials in the ideal. Also, the general approach of Ahern and Clark has been expanded upon in several different ways. The text of Chen and Guo [CG03] is an in-depth exploration with further results in this direction [[, see also]]DP89. For certain operator theoretic characterization of shift-invariant subspaces, see [MMSS19, Man88, SSW13].
Soltani gave an equivalent condition for shift-cyclicity of an outer function in $H^{2}(\mathbb{D}^{2})$ via stationary random fields [[, see]Theorem 2.18]Sol84, however, from a function theoretic point of view, it leaves a lot to be desired. For instance, it is not clear if the properties mentioned in Section 2.1.3 generalize to $d>1$. The following result of Ginsberg, Neuwirth and Newman extends SCP to $d>1$ [[, see]Theorem 5]GNN70666The original statement of Theorem 5 in [GNN70] is only valid for $p=2$ but the argument generalizes easily to all the other values of $p$..
@@
It is known that $H^{\infty}(\mathbb{D}^{d})$ is the multiplier algebra of $H^{p}(\mathbb{D}^{d})$ for $0<p<\infty$ [[, see]Section 2.1 (8)]Nik12, i.e.,
@@
Moreover, MC extends to $d>1$ [[, see]Section 2.1 (9)]Nik12, however, we do not know if NI, Mult or Inv generalize to $d>1$. The following result, which is a slightly modified version of a result of Nikolski [[, see]Theorem 3.3 (3)]Nik19, shows that Inv ‘almost’ holds when $d>1$.
@@
Since we do not know if NI holds for $d>1$, we cannot conclude that $f$ above is shift-cyclic in $H^{p}(\mathbb{D}^{d})$. In fact, if we can find an $f\in H^{p}(\mathbb{D}^{d})$ such that $1/f\in H^{q}(\mathbb{D}^{d})$ for some $0<p,q<\infty$ but $f$ is not shift-cyclic in $H^{p}(\mathbb{D}^{d})$, then we can conclude that NI and Mult do not generalize as such to $d>1$ [[, see]Proposition 2.1.1]Sam22. This brings us to some interesting open problems.
@@ 2.3. On the unit multidisk $\mathbb{D}^{\infty}_{2}$
Hilbert [Hil09] introduced a theory of holomorphic functions on the unit multidisk (also, Hilbert multidisk) defined as
$\mathbb{D}^{\infty}_{2}:=\left\{w=(w_{1},w_{2},\dots):|w_{n}|<1\text{ for all }n\in\mathbb{N}\text{ and }w\in\ell^{2}(\mathbb{C})\right\}.$
@@
It is not difficult to check that any power-series whose coefficients satisfy the given condition in (2.4) is holomorphic in the sense of [Hil09], and that it converges uniformly and absolutely on compact subsets of $\mathbb{D}^{\infty}_{2}$ [[, see]Lemma 6.6.2]Nik19. Now, $H^{2}(\mathbb{D}^{\infty}_{2})$ turns into a Hilbert space under the inner product borrowed from $\ell^{2}(\mathbb{Z}^{\infty}_{+})$ by the coefficients. Moreover, it can be checked that $H^{2}(\mathbb{D}^{\infty}_{2})$ satisfies appropriate modifications of P1-P3 [[, see]Lemma 6.6.2]Nik19.
@@
where $w\in\mathbb{D}^{\infty}_{2}$ is arbitrary [[, see]Corollary 3.7]Nik12.
Let $H^{\infty}(\mathbb{D}^{\infty}_{2})$ be the space of all bounded holomorphic functions on $\mathbb{D}^{\infty}_{2}$. It is possible to define $H^{p}(\mathbb{D}^{\infty}_{2})$ for any $0<p<\infty$ as well. The curious reader can find more information on this fact and the following sample of results about shift-cyclicity in $H^{2}(\mathbb{D}^{\infty}_{2})$ in [Nik12] (see also the correction from [Nik18]).
@@
Other interesting partial results along the same lines were obtained in [DG21]. For partial results on the generalization of Theorem 2.16, see [DGH18]. Let us now see how shift-cyclicity in $H^{2}(\mathbb{D}^{\infty}_{2})$ connects to, and, in fact, is equivalent to some other interesting problems as hinted in Example 1.5.
@@ 2.3.1. Function theory junction
The problem of determining $\vartheta$-cyclic functions is called the periodic dilation completeness problem (PDCP). For $\vartheta$ as in Example 1.5 $(3)$, determining when $\chi_{[0,1]}\in\vartheta[\varphi]$ for a given $\varphi\in L^{2}(0,\infty)$ is called the dilation completeness problem (DCP). Theorem 1.6 shows that solving DCP also solves the Riemann hypothesis, but $\varrho$ in the theorem is not $2$-periodic. The connection between PDCP and the Riemann hypothesis was observed by Noor [Noo19]. Wintner [Win44], Nyman [Nym50] and Beurling [Beu55] were the first to observe the relationship between PDCP and certain properties of Dirichlet series. In modern literature, the works of Hedenmalm, Lindquist and Seip [HLS97], Bagchi [Bag06], Nikolski [Nik12, Nik18, Nik19] and Noor [Noo19] have been highly influential. Let $H^{2}_{0}(\mathbb{D})$ be as in Example 1.5 $(2)$ and note that $L^{2}(0,1)$ and $H^{2}_{0}(\mathbb{D})$ are unitarily equivalent via
@@
Under the same correspondence, it is also clear that $\vartheta$-cyclicity is equivalent to $\tau$-cyclicity (again, from Example 1.5 $(2)$). Taking things one step further we can show that $\tau$-cyclicity is equivalent to shift-cyclicity in $H^{2}(\mathbb{D}^{\infty}_{2})$. Indeed, we can use an idea of Bohr [Boh13] for this. Let us order all the prime numbers $p_{m}$ in increasing order along $m\in\mathbb{N}$. Every $n\in\mathbb{N}$ has a unique prime factorization
@@
$\mathfrak{B}:\sum_{n\in\mathbb{N}}c_{n}z^{n}\in H^{2}_{0}(\mathbb{D})\mapsto\sum_{\alpha\in\mathbb{Z}_{+}^{d}}c_{\kappa^{-1}(\alpha)}z^{\alpha}\in H^{2}(\mathbb{D}^{\infty}_{2}).$
It is a standard exercise to now check that $\mathfrak{B}:H^{2}_{0}(\mathbb{D})\to H^{2}(\mathbb{D}^{\infty}_{2})$ is a unitary and $\tau$-cyclicity is equivalent to shift-cyclicity in $H^{2}(\mathbb{D}^{\infty}_{2})$ [[, see]Section 6.6.3]Nik19.
While things are clear from an operator theoretic point of view, notice that point-wise relationships are lost under these unitary equivalences. For instance, Nikolski [Nik12] used a theorem of Green and Tao [GT08] about the existence of arbitrarily long arithmetic progressions among the prime numbers to show that
@@
[[, see also]Exercise 6.7.4 (c)]Nik19. Moreover, from P2 it is clear that if $F$ is shift-cyclic in $H^{2}(\mathbb{D}^{\infty}_{2})$ then it is non-vanishing on $\mathbb{D}^{\infty}_{2}$, but it is not clear what this property corresponds to under $\mathfrak{B}^{-1}$. We therefore end this subsection with the following open problem.
@@ 3.1. On the unit disk $\mathbb{D}$
In this section, we present a situation where shift-cyclicity is mysterious despite functions in the space exhibiting the same factorization and boundary properties as the one variable Hardy spaces, but the reason this time is partly due to functions having a bit more rigidity along the boundary. We follow the notation from [EKMR14] throughout this section.

### v3

@@ 1. SCP
    Note that $f(z)=z-w$ satisfies (2.2) exactly when $w\not\in\mathbb{D}$ (see [94, Lemma 15.17]). If we combine Theorem 2.10 and Proposition 1.12, we get that $P$ is shift-cyclic in any of the Hardy spaces if and only if it is non-vanishing on $\mathbb{D}$. For a proof that does not use Theorem 2.10, see [41, Theorem 5].
@@
    For any $f\in H^{p}(\mathbb{D})$ with $0<p<\infty$, a result of Garnett [39, Chapter II Theorem 7.4] then shows that
@@ 5. Inv
SCP refers to the shift-cyclicity of polynomials. It will be a general theme moving forward that the vanishing set of a polynomial plays a crucial role in determining its shift-cyclicity.
@@
Multiplier-cyclicity is a natural substitute for shift-cyclicity in non-analytic function spaces. We shall learn more about this topic in Section 4.2.
@@ 2.2. On the unit polydisk $\mathbb{D}^{d}$
The boundary properties of $H^{p}(\mathbb{D})$ functions are explained by the results of Fatou [37] and Riesz [90]. These results are, in essence, consequences of the symmetries of the Poisson kernel and the Poisson integral representation of Hardy space functions. Many of these properties can be generalized to $d>1$ via slice functions, i.e., given any $f\in\operatorname{Hol}(\mathbb{D}^{d})$ and $\zeta\in\mathbb{T}^{d}$, consider
@@
exist a.e. on $\mathbb{T}^{d}$ [93, Theorem 3.3.3]. Bochner [21] used slice functions to extend Riesz’s theorem and showed that $\log|f^{*}|\in L^{1}(\mathbb{T}^{d})$ unless $f\equiv 0$, and
@@
for all $f\in H^{p}(\mathbb{D}^{d})$ with $0<p<\infty$ (see [93, Theorem 3.4.3] for a different treatment). It is then not difficult to verify that Corollary 2.6 and Theorem 2.7 can be generalized for $d>1$. Unfortunately, we cannot go too far from here in terms of factorization and shift-cyclicity when $d>1$.
@@ 2.2.1. Negative results
All examples mentioned in this subsection are non-trivial, as well as technical, so we omit the details to keep our discussion compact. The reader is encouraged to look at the references below for these details. Let us start with the following result of Rudin (see [93, corollary to Theorem 4.4.2]) which shows that Theorem 2.10 cannot be generalized as is to the case $d>1$.
@@
One might wonder if we could provide an appropriate description of those shift-invariant subspaces that are finitely generated. The following result of Miles [76] sheds some light on this. Similar to the one variable case, we let $\mathcal{Z}(f)$ denote the zero set of any $f\in\operatorname{Hol}(\mathbb{D}^{d})$ in $\mathbb{D}^{d}$ counting multiplicities [^5] The multiplicity of a zero at $w$ is the smallest integer $k$ such that the $k$-th term in the homogeneous expansion of $f$ around $w$ does not vanish at $w$ (see [93, Section 1.1])..
@@
Combined with Theorem 2.12, the above inclusion gives the following result. Note that while Miles’ result holds when $d=2$, one can easily generalize it to any $d>1$ by taking
@@
The penultimate nail in the coffin is that there is no real hope for a useful factorization result for $d>1$ either. Firstly, Theorem 2.12 shows that the result of F. and R. Nevanlinna [79] (see p. 2.3) does not hold for $d>1$. Next, we have the following result of Rosay [91] (generalizing a result of Rudin [92] and Miles [77]). Recall from Mult that the Smirnov factorization result implies that every $H^{1}(\mathbb{D})$ function can be factored into a product of two $H^{2}(\mathbb{D})$ functions.
@@
Finally, if one only cares about shift-cyclicity of a function, then it is natural to ask if there is an analogue of an outer function when $d>1$. Rudin defines an outer function as a function $f\in H^{p}(\mathbb{D}^{d})$ such that $f(0)\neq 0$ and
$\log|f(0)|=\int_{\mathbb{T}^{d}}\log|f^{*}|,$
@@
In [93, Theorem 4.4.6], Rudin shows that every shift-cyclic function in $H^{p}(\mathbb{D}^{d})$ is outer. Thus, one-half of the Smirnov–Beurling holds. However, in [93, Theorem 4.4.8 (b)], Rudin then constructs an explicit example of an outer function in $H^{2}(\mathbb{D}^{2})$ that is not shift-cyclic.
@@ 2.2.2. Positive results
We start with a result of Ahern and Clark which shows that all shift-invariant subspaces of $H^{2}(\mathbb{D}^{d})$ with finite codimension are generated by finitely many polynomials (see [3, Theorem 3 and Corollary 4]).
@@
Finite codimensionality of polynomial ideals is fairly well-understood, both algebraically and computationally (see [65, Section 3.7]). In fact, the codimension of such an ideal is precisely the number of common zeros (in $\mathbb{D}^{d}$) of polynomials in the ideal. Also, the general approach of Ahern and Clark has been expanded upon in several different ways. The text of Chen and Guo [26] is an in-depth exploration with further results in this direction (also see [32]). For certain operator theoretic characterization of shift-invariant subspaces, see [70, 71, 98].
Soltani gave an equivalent condition for shift-cyclicity of an outer function in $H^{2}(\mathbb{D}^{2})$ via stationary random fields (see [102, Theorem 2.18]), however, from a function theoretic point of view, it leaves a lot to be desired. For instance, it is not clear if the properties mentioned in Section 2.1.3 generalize to $d>1$. The following result of Ginsberg, Neuwirth and Newman extends SCP to $d>1$ (see [41, Theorem 5]) [^6] The original statement of Theorem 5 in [41] is only valid for $p=2$ but the argument generalizes easily to all the other values of $p$..
@@
It is known that $H^{\infty}(\mathbb{D}^{d})$ is the multiplier algebra of $H^{p}(\mathbb{D}^{d})$ for $0<p<\infty$ (see [81, Section 2.1 (8)]), i.e.,
@@
Moreover, MC extends to $d>1$ (see [81, Section 2.1 (9)]), however, we do not know if NI, Mult or Inv generalize to $d>1$. The following result, which is a slightly modified version of a result of Nikolski (see [81, Theorem 3.3 (3)]), shows that Inv ‘almost’ holds when $d>1$.
@@
Since we do not know if NI holds for $d>1$, we cannot conclude that $f$ above is shift-cyclic in $H^{p}(\mathbb{D}^{d})$. In fact, if we can find an $f\in H^{p}(\mathbb{D}^{d})$ such that $1/f\in H^{q}(\mathbb{D}^{d})$ for some $0<p,q<\infty$ but $f$ is not shift-cyclic in $H^{p}(\mathbb{D}^{d})$, then we can conclude that NI and Mult do not generalize as such to $d>1$ (see [97, Proposition 2.1.1]). This brings us to some interesting open problems.
@@ 2.3. On the unit multidisk $\mathbb{D}^{\infty}_{2}$
Hilbert [51] introduced a theory of holomorphic functions on the unit multidisk (also, Hilbert multidisk) defined as
$\mathbb{D}^{\infty}_{2}:=\left\{w=(w_{1},w_{2},\dots):|w_{n}|<1\text{ for all }n\in\mathbb{N}\text{ and }w\in\ell^{2}(\mathbb{N})\right\}.$
@@
It is not difficult to check that any power-series whose coefficients satisfy the given condition in (2.4) is holomorphic in the sense of [51], and that it converges uniformly and absolutely on compact subsets of $\mathbb{D}^{\infty}_{2}$ (see [83, Lemma 6.6.2]). Now, $H^{2}(\mathbb{D}^{\infty}_{2})$ turns into a Hilbert space under the inner product borrowed from $\ell^{2}(\mathbb{Z}^{\infty}_{+})$ by the coefficients. Moreover, it can be checked that $H^{2}(\mathbb{D}^{\infty}_{2})$ satisfies appropriate modifications of P1-P3 (see [83, Lemma 6.6.2]).
@@
where $w\in\mathbb{D}^{\infty}_{2}$ is arbitrary (see [81, Corollary 3.7]).
Let $H^{\infty}(\mathbb{D}^{\infty}_{2})$ be the space of all bounded holomorphic functions on $\mathbb{D}^{\infty}_{2}$. It is possible to define $H^{p}(\mathbb{D}^{\infty}_{2})$ for any $0<p<\infty$ as well. The curious reader can find more information on this fact and the following sample of results about shift-cyclicity in $H^{2}(\mathbb{D}^{\infty}_{2})$ in [81] (see also the correction from [82]).
@@
Other interesting partial results along the same lines were obtained in [29]. For partial results on the generalization of Theorem 2.16, see [30]. Let us now see how shift-cyclicity in $H^{2}(\mathbb{D}^{\infty}_{2})$ connects to, and, in fact, is equivalent to some other interesting problems as hinted in Example 1.5.
@@ 2.3.1. Function theory junction
The problem of determining $\vartheta$-cyclic functions is called the periodic dilation completeness problem (PDCP). For $\vartheta$ as in Example 1.5 $(3)$, determining when $\chi_{[0,1]}\in\vartheta[\varphi]$ for a given $\varphi\in L^{2}(0,\infty)$ is called the dilation completeness problem (DCP). Theorem 1.6 shows that solving DCP also solves the Riemann hypothesis, but $\varrho$ in the theorem is not $2$-periodic. The connection between PDCP and the Riemann hypothesis was observed by Noor [84]. Wintner [106], Nyman [85] and Beurling [20] were the first to observe the relationship between PDCP and certain properties of Dirichlet series. In modern literature, the works of Hedenmalm, Lindquist and Seip [50], Bagchi [11], Nikolski [81, 82, 83] and Noor [84] have been highly influential.
Let $H^{2}_{0}(\mathbb{D})$ be as in Example 1.5 $(2)$ and note that $L^{2}(0,1)$ and $H^{2}_{0}(\mathbb{D})$ are unitarily equivalent via
@@
Under the same correspondence, it is also clear that $\vartheta$-cyclicity is equivalent to $\tau$-cyclicity (again, from Example 1.5 $(2)$). Taking things one step further we can show that $\tau$-cyclicity is equivalent to shift-cyclicity in $H^{2}(\mathbb{D}^{\infty}_{2})$. Indeed, we can use an idea of Bohr [22] for this. Let us order all the prime numbers $p_{m}$ in increasing order along $m\in\mathbb{N}$. Every $n\in\mathbb{N}$ has a unique prime factorization
@@
$\mathfrak{B}:\sum_{n\in\mathbb{N}}c_{n}z^{n}\in H^{2}_{0}(\mathbb{D})\mapsto\sum_{\alpha\in\mathbb{Z}_{+}^{\infty}}c_{\kappa^{-1}(\alpha)}z^{\alpha}\in H^{2}(\mathbb{D}^{\infty}_{2}).$
It is a standard exercise to check that $\mathfrak{B}:H^{2}_{0}(\mathbb{D})\to H^{2}(\mathbb{D}^{\infty}_{2})$ is a unitary and $\tau$-cyclicity is equivalent to shift-cyclicity in $H^{2}(\mathbb{D}^{\infty}_{2})$ (see [83, Section 6.6.3]).
While things are clear from an operator theoretic point of view, notice that point-wise relationships are lost under these unitary equivalences. For instance, Nikolski [81] used a theorem of Green and Tao [43] about the existence of arbitrarily long arithmetic progressions among the prime numbers to show that
@@
(also see [83, Exercise 6.7.4 (c)]). Moreover, from P2 it is clear that if $F$ is shift-cyclic in $H^{2}(\mathbb{D}^{\infty}_{2})$ then it is non-vanishing on $\mathbb{D}^{\infty}_{2}$, but it is not clear what this property corresponds to under $\mathfrak{B}^{-1}$. We therefore end this subsection with the following open problem.
@@ 3.1. On the unit disk $\mathbb{D}$
In this section, we present a situation where shift-cyclicity is mysterious despite functions in the space exhibiting the same factorization and boundary properties as the one variable Hardy spaces, but the reason this time is partly due to functions having a bit more rigidity along the boundary. We follow the notation from [35] throughout this section.